OpenStudy (anonymous):

I need some help pls. The viera family is traveling from NY to Orlando for vacation the equation y=1000-65xrepresents the distance remaining in their trip after x hours. A) graph the equation to find the distance remaining after 6hrs. B) interpret the slope and the y intercept.

OpenStudy (anonymous):

for part (A), you just need to graph this line... you can put in different x values and see what y values result. So after x = 1 hour, y will be y = 1000 - 65(1) = 935, and so on for x = 2, x = 3, x = 4, x= 5, and finally x = 6. Part A asks you to find the value of y... the distance remaining... when x = 6 hours. But it sounds like you are supposed to graph it. For part (B), reorganize the equation into the normal "slope-intercept" form, like y = mx+b. y = 1000 - 65x y = -65x + 1000 y = mx + b Then remember that slope is m and y-intercept is b.

OpenStudy (anonymous):

thanks very much Jake v8

OpenStudy (anonymous):

you are welcome... glad to help :)

OpenStudy (anonymous):

can some one please help me with this problem please...If y=-12 when x=9,find y when x=-4...i got 5.3repeating